Use the data in the table below to determine the rate constant (in L mol-1s-1) for the reaction below? 2?2(?)+?2(?)⇌2?2?(?) 2 H 2 ( g ) + O 2 ( g ) ⇌ 2 H 2 O ( g ) Experiment [H2]0 (M) [O2]0 (M) Initial rate (M s-1) 1 2.5 x 10-2 1.0 x 10-2 4 2 5.0 x 10-2 1.0 x 10-2 8 3 2.5 x 10-2 0.5 x 10-2 2

Select one:

a. 1.2 x 103 3

b. 1.2 x 1011 11

c. 1.6 x 104 4

d. 4.0 x 1011 11

e. 6.0 x 103 3
Answer #1

see experiment 1 and 2:

[H2] doubles

[O2] is constant

rate doubles

so, order of H2 is 1

see experiment 3 and 1:

[H2] is constant

[O2] doubles

rate doubles

so, order of O2 is 1

overall order = 1 + 1 = 2

Rate law is:

rate = k*[H2]*[O2]

Put values from 1st row of table in rate law

rate = k*[H2]*[O2]

4 = k*0.025*0.01

k = 1.6*10^4 M-1.s-1

Answer: c